# ACT-Tensor: Tensor Completion Framework for Financial Dataset Imputation

### v1

@@ 3.2.2 Cluster-based Completion
Dense Clusters. For dense clusters, we apply tensor completion directly to their own sub-tensors. For each $d\in\mathcal{D}$, the sub-tensor of dense cluster $d$ is represnted as:
$\mathcal{X}_{d}=\mathcal{X}_{:\,,\,\mathcal{I}_{d}\,,\,:}\;\in\;\mathbb{R}^{T\times|\mathcal{I}_{d}|\times L}$ (8)
[…]
$\min_{U,\;V_{d},\;W}\bigl{\|}\mathcal{P}_{\Omega_{d}}(\mathcal{X}_{d})-\mathcal{P}_{\Omega_{d}}\left([\![U,V_{d},W]\!]\right)\bigr{\|}_{F}^{2},$ (9)
@@
$\mathcal{X}^{\rm agg}_{s}=\mathcal{X}_{:\,,\,\mathcal{I}_{s}^{\rm agg}\,,\,:}\in\mathbb{R}^{T\times|\mathcal{I}_{s}^{\rm agg}|\times L},\quad\text{with}\;\mathcal{I}_{s}^{\rm agg}=\mathcal{I}_{s}\cup\bigcup_{d\in\mathcal{D}}\mathcal{I}_{d}.$ (10)
[…]
$\begin{split}\min_{U,\,V^{\rm agg}_{s},\,W}\;&\left\|\mathcal{P}_{\Omega^{\rm agg}_{s}}(\mathcal{X}^{\rm agg}_{s})-\mathcal{P}_{\Omega^{\rm agg}_{s}}\left([\![U,V^{\rm agg}_{s},W]\!]\right)\right\|_{F}^{2}\\ &\quad+\lambda\left(\|U\|_{F}^{2}+\|V^{\rm agg}_{s}\|_{F}^{2}+\|W\|_{F}^{2}\right).\end{split}$ (11)
@@
$\hat{\mathcal{X}}_{d}=(\hat{\mathcal{X}}^{\mathrm{agg}}_{s})_{:\,,\,\mathcal{I}_{s}\,,\,:}\;\in\;\mathbb{R}^{T\times|\mathcal{I}_{s}|\times L}.$ (12)
Repeating this for every $s\in\mathcal{S}$ produces completed sub-tensors for all sparse clusters.
3.2.3 Assembly of Completed Sub-tensors
@@
$\hat{\mathcal{X}}_{:,\mathcal{I}_{k},:}=\hat{\mathcal{X}}_{k},\quad\text{with}\;k=1,\dots,K.$ (13)
Concatenating all $K$ completed sub-tensors yields the fully imputed tensor denoted by $\hat{\mathcal{X}}$, which is then passed to the temporal-smoothing module.
3.3 Temporal Smoothing
While cluster-aware completion captures high-level time patterns through its temporal factors, it can still leave short-lived noise that obscures the true sequential dynamics. To restore underlying dynamics, we add a temporal smoothing module that filters each firm–characteristic series before the final panel is used in asset-pricing tests. We consider three smoother options: centered moving average (Section 3.3.1), exponential moving average (Section 3.3.2), or Kalman filter (Section 3.3.3).
3.3.1 Centered Moving Average (CMA)
To smooth out short-lived noise without flattening medium-term trends in the data, we apply a centered moving average to each imputed series. Specifically, for window size $\delta$ with $m=(\delta-1)/2$:
$\tilde{x}_{t}^{n,\ell}=\frac{1}{\delta}\sum_{s=-m}^{m}\hat{x}_{t+s}^{n,\ell}\quad\text{with}\;t=m+1,\dots,T-m$ (14)
where $\hat{x}_{t+s}^{n,\ell}$ denotes the imputed data of company $n$’s characteristic $\ell$ at $t+s$. Near the ends of the series ($t\leq m$ or $t>T-m$), we shrink the window to include only available observations.
3.3.2 Exponential Moving Average (EMA)
@@
$\tilde{x}_{1}^{n,\ell}=\hat{x}_{1}^{n,\ell},\quad\tilde{x}_{t}^{n,\ell}=\theta\,\hat{x}_{t}^{n,\ell}+(1-\theta)\,\tilde{x}_{t-1}^{n,\ell},\quad\text{with}\;t=2,\dots,T.$ (15)
3.3.3 Kalman Filter (KF)
@@
$y_{t}^{n,\ell}=y_{t-1}^{n,\ell}+w_{t},\quad\text{with}\;w_{t}\sim\mathcal{N}(0,h),$ (16)
@@
$\hat{x}_{t}^{n,\ell}=y_{t}^{n,\ell}+v_{t},\quad\text{with}\;v_{t}\sim\mathcal{N}(0,r).$ (17)
where $h$ and $r$ controls process- and measurement-noise variance, respectively. Starting from a prior $(y_{1}^{n,\ell},P_{1}^{n,\ell})$, the standard predict–update recursion produces filtered means $\bar{y}_{t}^{n,\ell}$ and variances $P_{t}^{n,\ell}$. A backward smoothing pass then yields
$\tilde{x}_{t}^{n,\ell}=\mathbb{E}\left[y_{t}^{n,\ell}\mid\hat{x}_{1:T}^{n,\ell}\right],$ (18)
@@
The full architecture of ClusterTensor is shown in Algorithm 1.
3.4 Imputation Accuracy Metrics
To assess how closely each method reconstructs the held-out characteristic values, we report four widely used error measures. Let the evaluation set contain $M$ masked entries for evaluation with true values $x_{m}$ and imputed values $\tilde{x}_{m}$ for $m=1,\ldots,M$:
$\text{RMSE}_{\text{imp}}$ $=\sqrt{\frac{1}{M}\sum_{m=1}^{M}(x_{m}-\tilde{x}_{m})^{2}},\quad\text{MAE}_{\text{imp}}=\frac{1}{M}\sum_{m=1}^{M}\lvert x_{m}-\tilde{x}_{m}\rvert,$ (19)
@@ 19 $\tilde{x}_{t}^{n,\ell}\leftarrow\mathrm{Smooth}\left(\hat{x}_{t}^{n,\ell};\,S,\,\alpha,\beta\right)$
4 Downstream Asset Pricing Evaluation
This section explores whether the imputed characteristic panels capture genuine financial signals rather than just noise with an advanced asset-pricing pipeline tailored for tensor-structured financial data [26]. We first convert the panel into investable strategies by forming value-weighted portfolios, creating an excess return tensor (Section 4.1). We compress this tensor into a small set of latent return drivers with tensor decomposition to identify which of those factors genuinely forecast future returns (Section 4.2). Then, we quantify their predictive strength and rebuild a filtered excess return tensor, providing the foundation for evaluation metrics reported later (Section 4.3).
4.1 Excess Return Tensor Construction
@@
$\mathcal{R}_{p,q,\ell,t}=\frac{\sum_{i\in\mathcal{B}_{p,q,\ell,t}}w_{i,t}\,r_{i,t}}{\sum_{i\in\mathcal{B}_{p,q,\ell,t}}w_{i,t}}-r_{f,t},$ (20)
where $w_{i,t}$ is firm $i$’s market capitalization, $r_{i,t}$ is firm $i$’s realized return, and $r_{f,t}$ is the risk-free rate. Stacking these returns over time and characteristics produces the portfolio-return tensor:
$\mathcal{R}\in\mathbb{R}^{P\times Q\times(L-1)\times T}.$ (21)
4.2 Tensor-Based Factor Extraction
@@
$\operatorname*{argmin}\|\mathcal{R}-{\mathcal{F}}\times_{2}U\times_{3}V\times_{4}W\|_{F}.$ (22)
As Equation 22 does not have a closed-form solution, we apply Higher-Order SVD [36] to approximate $U$, $V$, $W$ and get $\hat{U}\in\mathbb{R}^{P\times k_{p}}$, $\hat{V}\in\mathbb{R}^{Q\times k_{q}}$, and $\hat{W}\in\mathbb{R}^{(L-1)\times k_{\ell}}$. After obtaining the loading matrices, we project back to get $\hat{\mathcal{F}}$:
^=×$\hat{\mathcal{F}}=\mathcal{R}\times_{2}\hat{U}^{\intercal}\times_{3}\hat{V}^{\intercal}\times_{4}\hat{W}^{\intercal},$ (23)
where ^$\hat{\mathcal{F}}\in\mathbb{R}^{k_{p}\times k_{q}\times k_{\ell}\times T}$. By stacking the $k=k_{p}\,k_{q}\,k_{\ell}$ time-mode slices of $\hat{\mathcal{F}}$ at each $t$ into $\mathbf{f}_{t}\in\mathbb{R}^{k}$, we obtain a compact latent state that is a concise summary of the key drivers of cross-sectional returns at $t$. This latent state serves as the input for the return-prediction regression described below.
4.3 Factor-Based Return Prediction
@@
${r}_{n,t+1}=\alpha_{n}+\boldsymbol{\beta}_{n}^{\intercal}\mathbf{f}_{t}^{(\mathcal{M})}+\varepsilon_{n,t+1},$ (24)
where $r_{n,t+1}$ is the excess return of portfolio $n$ next period, $\mathbf{f}_{t}^{(\mathcal{M})}$ is a candidate subset of the $k$ latent factors, and $\alpha_{n}$ is the pricing-error intercept. We choose the subset $\mathcal{M}$ with a forward stepwise search that maximize the pseudo cross-sectional $R^{2}$:
$R^{2}_{\text{xs}}=1-\frac{\frac{1}{N}\sum_{n=1}^{N}\alpha_{n}^{2}}{\displaystyle\operatorname{Var}_{\text{xs}}(\bar{R}_{n})},$ (25)
adding one factor at a time until the target model size $|\mathcal{M}|$ is reached. This yields the most informative, yet parsimonious, factor set for prediction. With the selected factor set and the estimated loadings, we can compute the model-implied forecasts:
$\hat{r}_{n,t+1}=\hat{\alpha}_{n}+\hat{\boldsymbol{\beta}}_{n}^{\intercal}\mathbf{f}_{t}^{(\mathcal{M})},$ (26)
@@
4.4 Asset Pricing Evaluation Metrics
The factor-filtered return tensor is assessed along two complementary dimensions: pricing accuracy and predictive power.
4.4.1 Pricing Accuracy.
For each portfolio $n$, the return-forecasting regression (Equation 26) yields a pricing error intercept $\alpha_{n}$. We summarize these errors with root-mean-squared and mean-absolute measure [26]:
${\text{RMSE}_{\alpha}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}\hat{\alpha}_{n}^{2}}},\qquad{\text{MAE}_{\alpha}=\frac{1}{N}\sum_{n=1}^{N}\left|\hat{\alpha}_{n}\right|}.$ (27)
Lower values indicate that the selected factors account for a larger share of the cross-sectional return variation.
4.4.2 Predictive Power.
@@ 4.4 Asset Pricing Evaluation Metrics
Following [9], we measure the gap between predicted and realized cross-sectional rankings with the mean-absolute rank error:
$\text{MAE\mbox{-}Rank}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{N}\sum_{n=1}^{N}\left|\operatorname{rank}\left(r_{n,t+1}\right)-\operatorname{rank}\left(\hat{r}_{n,t+1}\right)\right|.$ (28)
A smaller MAE-Rank value indicates fewer mis-ordered portfolios.
The Information Coefficient (IC) [37] captures the linear association between predictions and outcomes:
$\text{IC}=\frac{1}{T}\sum_{t=1}^{T}\operatorname{corr}\left(\hat{\mathbf{r}}_{t+1},\,\mathbf{r}_{t+1}\right),$ (29)
@@
To turn ranking skill into investable profits, each month we form a Top-minus-Bottom (T-B) portfolio by longing the decile of firms with the highest predicted excess returns $\hat{r}_{n,t+1}$ and shorting the decile with the lowest [2, 3]. The resulting excess return is:
$r^{\text{T-B}}_{t+1}=\frac{1}{N_{\text{top}}}\sum_{i\in\text{Top}}r_{i,t+1}-\frac{1}{N_{\text{bot}}}\sum_{i\in\text{Bottom}}r_{i,t+1}.$ (30)
where $N_{\text{top}}$ and $N_{\text{bot}}$ are the numbers of firms in the two deciles. We summarize the strategy’s risk-adjusted performance with its annualized Sharpe ratio (T-B Sharpe):
$\text{Sharpe}_{\text{T-B}}=\sqrt{12}\cdot\frac{\bar{r}^{\text{T-B}}}{\sigma^{\text{T-B}}}.$ (31)
@@
5 Experiment
To evaluate the effectiveness of our imputation framework, we adopt a two-step evaluation that assesses both the statistical accuracy of the data imputation and the financial utility that the imputed data provides in asset-pricing applications.

### v3

@@ 3.2.2. Cluster-based Completion
Dense Clusters. For dense clusters, we apply tensor completion directly to their own sub-tensors. For each $d\in\mathcal{D}$, the sub‐tensor of dense cluster $d$ is represnted as:
(8) $\mathcal{X}_{d}=\mathcal{X}_{:\,,\,\mathcal{I}_{d}\,,\,:}\;\in\;\mathbb{R}^{T\times|\mathcal{I}_{d}|\times L}$
We then solve a rank-$R$ CP problem on its observed entries, setting $\lambda=0$ because over‑fitting is not a concern at this density:
(9) $\min_{U,\;V_{d},\;W}\bigl\|\mathcal{P}_{\Omega_{d}}(\mathcal{X}_{d})-\mathcal{P}_{\Omega_{d}}\left([\![U,V_{d},W]\!]\right)\bigr\|_{F}^{2},$
@@
(10) $\mathcal{X}^{\rm agg}_{s}=\mathcal{X}_{:\,,\,\mathcal{I}_{s}^{\rm agg}\,,\,:}\in\mathbb{R}^{T\times|\mathcal{I}_{s}^{\rm agg}|\times L},\quad\text{with}\;\mathcal{I}_{s}^{\rm agg}=\mathcal{I}_{s}\cup\bigcup_{d\in\mathcal{D}}\mathcal{I}_{d}.$
We then solve a rank-$R$ CP problem on the observed entries of this aggregated tensor, using a $\ell_{2}$ penalty:
(11) $\begin{split}\min_{U,\,V^{\rm agg}_{s},\,W}\;&\left\|\mathcal{P}_{\Omega^{\rm agg}_{s}}(\mathcal{X}^{\rm agg}_{s})-\mathcal{P}_{\Omega^{\rm agg}_{s}}\left([\![U,V^{\rm agg}_{s},W]\!]\right)\right\|_{F}^{2}\\ &\quad+\lambda\left(\|U\|_{F}^{2}+\|V^{\rm agg}_{s}\|_{F}^{2}+\|W\|_{F}^{2}\right).\end{split}$
@@
(12) $\hat{\mathcal{X}}_{d}=(\hat{\mathcal{X}}^{\mathrm{agg}}_{s})_{:\,,\,\mathcal{I}_{s}\,,\,:}\;\in\;\mathbb{R}^{T\times|\mathcal{I}_{s}|\times L}.$
Repeating this for every $s\in\mathcal{S}$ produces completed sub‐tensors for all sparse clusters.
3.2.3. Assembly of Completed Sub-tensors
@@
(13) $\hat{\mathcal{X}}_{:,\mathcal{I}_{k},:}=\hat{\mathcal{X}}_{k},\quad\text{with}\;k=1,\dots,K.$
Concatenating all $K$ completed sub-tensors yields the fully imputed tensor denoted by $\hat{\mathcal{X}}$, which is then passed to the temporal‐smoothing module.
3.3. Temporal Smoothing
While cluster‑aware completion captures high-level time patterns through its temporal factors, it can still leave short‑lived noise that obscures the true sequential dynamics. To restore underlying dynamics, we add a temporal smoothing module that filters each firm–characteristic series before the final panel is used in asset‑pricing tests. We consider three smoother options: centered moving average (Section 3.3.1), exponential moving average (Section 3.3.2), or Kalman filter (Section 3.3.3).
3.3.1. Centered Moving Average (CMA)
To smooth out short-lived noise without flattening medium‑term trends in the data, we apply a centered moving average to each imputed series. Specifically, for window size $\delta$ with $m=(\delta-1)/2$:
(14) $\tilde{x}_{t}^{n,\ell}=\frac{1}{\delta}\sum_{s=-m}^{m}\hat{x}_{t+s}^{n,\ell}\quad\text{with}\;t=m+1,\dots,T-m$
where $\hat{x}_{t+s}^{n,\ell}$ denotes the imputed data of company $n$’s characteristic $\ell$ at $t+s$. Near the ends of the series ($t\leq m$ or t>Tmt>T-m), we shrink the window to include only available observations.
3.3.2. Exponential Moving Average (EMA)
@@
(15) $\tilde{x}_{1}^{n,\ell}=\hat{x}_{1}^{n,\ell},\quad\tilde{x}_{t}^{n,\ell}=\theta\,\hat{x}_{t}^{n,\ell}+(1-\theta)\,\tilde{x}_{t-1}^{n,\ell},\quad\text{with}\;t=2,\dots,T.$
3.3.3. Kalman Filter (KF)
@@
(16) $y_{t}^{n,\ell}=y_{t-1}^{n,\ell}+w_{t},\quad\text{with}\;w_{t}\sim\mathcal{N}(0,h),$
@@
(17) $\hat{x}_{t}^{n,\ell}=y_{t}^{n,\ell}+v_{t},\quad\text{with}\;v_{t}\sim\mathcal{N}(0,r).$
where $h$ and $r$ controls process‐ and measurement-noise variance, respectively. Starting from a prior $(y_{1}^{n,\ell},P_{1}^{n,\ell})$, the standard predict–update recursion produces filtered means $\bar{y}_{t}^{n,\ell}$ and variances $P_{t}^{n,\ell}$. A backward smoothing pass then yields
(18) $\tilde{x}_{t}^{n,\ell}=\mathbb{E}\left[y_{t}^{n,\ell}\mid\hat{x}_{1:T}^{n,\ell}\right],$
@@
3.4. Imputation Accuracy Metrics
To assess how closely each method reconstructs the held‑out characteristic values, we report four widely used error measures. Let the evaluation set contain $M$ masked entries for evaluation with true values $x_{m}$ and imputed values $\tilde{x}_{m}$ for $m=1,\ldots,M$:
(19) $\text{RMSE}_{\text{imp}}$ $=\sqrt{\frac{1}{M}\sum_{m=1}^{M}(x_{m}-\tilde{x}_{m})^{2}},\quad\text{MAE}_{\text{imp}}=\frac{1}{M}\sum_{m=1}^{M}\lvert x_{m}-\tilde{x}_{m}\rvert$,
@@ 19 $\tilde{x}_{t}^{n,\ell}\leftarrow\mathrm{Smooth}\left(\hat{x}_{t}^{n,\ell};\,S,\,\alpha,\beta\right)$
4. Downstream Asset Pricing Evaluation
This section explores whether the imputed characteristic panels capture genuine financial signals rather than just noise with an advanced asset‐pricing pipeline tailored for tensor-structured financial data (Lettau, 2024). We first convert the panel into investable strategies by forming value‑weighted portfolios, creating an excess return tensor (Section 4.1). We compress this tensor into a small set of latent return drivers with tensor decomposition to identify which of those factors genuinely forecast future returns (Section 4.2). Then, we quantify their predictive strength and rebuild a filtered excess return tensor, providing the foundation for evaluation metrics reported later (Section 4.3).
4.1. Excess Return Tensor Construction
@@
(20) $\mathcal{R}_{p,q,\ell,t}=\frac{\sum_{i\in\mathcal{B}_{p,q,\ell,t}}w_{i,t}\,r_{i,t}}{\sum_{i\in\mathcal{B}_{p,q,\ell,t}}w_{i,t}}-r_{f,t},$
where $w_{i,t}$ is firm $i$’s market capitalization, $r_{i,t}$ is firm $i$’s realized return, and $r_{f,t}$ is the risk-free rate. Stacking these returns over time and characteristics produces the portfolio‑return tensor:
(21) $\mathcal{R}\in\mathbb{R}^{P\times Q\times(L-1)\times T}.$
4.2. Tensor‑Based Factor Extraction
@@
(22) $\operatorname*{argmin}\|\mathcal{R}-{\mathcal{F}}\times_{2}U\times_{3}V\times_{4}W\|_{F}.$
As Equation 22 does not have a closed-form solution, we apply Higher-Order SVD (De Lathauwer et al., 2000) to approximate $U$, $V$, $W$ and get $\hat{U}\in\mathbb{R}^{P\times k_{p}}$, $\hat{V}\in\mathbb{R}^{Q\times k_{q}}$, and $\hat{W}\in\mathbb{R}^{(L-1)\times k_{\ell}}$. After obtaining the loading matrices, we project back to get $\hat{\mathcal{F}}$:
(23) $\hat{\mathcal{F}}=\mathcal{R}\times_{2}\hat{U}^{\intercal}\times_{3}\hat{V}^{\intercal}\times_{4}\hat{W}^{\intercal},$
where $\hat{\mathcal{F}}\in\mathbb{R}^{k_{p}\times k_{q}\times k_{\ell}\times T}$. By stacking the $k=k_{p}\,k_{q}\,k_{\ell}$ time‐mode slices of $\hat{\mathcal{F}}$ at each $t$ into $\mathbf{f}_{t}\in\mathbb{R}^{k}$, we obtain a compact latent state that is a concise summary of the key drivers of cross-sectional returns at $t$. This latent state serves as the input for the return-prediction regression described below.
4.3. Factor-Based Return Prediction
@@
(24) ${r}_{n,t+1}=\alpha_{n}+\boldsymbol{\beta}_{n}^{\intercal}\mathbf{f}_{t}^{(\mathcal{M})}+\varepsilon_{n,t+1},$
where $r_{n,t+1}$ is the excess return of portfolio $n$ next period, $\mathbf{f}_{t}^{(\mathcal{M})}$ is a candidate subset of the $k$ latent factors, and $\alpha_{n}$ is the pricing‑error intercept. We choose the subset $\mathcal{M}$ with a forward stepwise search that maximize the pseudo cross-sectional $R^{2}$:
(25) $R^{2}_{\text{xs}}=1-\frac{\frac{1}{N}\sum_{n=1}^{N}\alpha_{n}^{2}}{\displaystyle\operatorname{Var}_{\text{xs}}(\bar{R}_{n})},$
adding one factor at a time until the target model size $|\mathcal{M}|$ is reached. This yields the most informative, yet parsimonious, factor set for prediction. With the selected factor set and the estimated loadings, we can compute the model‑implied forecasts:
(26) $\hat{r}_{n,t+1}=\hat{\alpha}_{n}+\hat{\boldsymbol{\beta}}_{n}^{\intercal}\mathbf{f}_{t}^{(\mathcal{M})},$
@@
4.4. Asset Pricing Evaluation Metrics
The factor‑filtered return tensor is assessed along two complementary dimensions: pricing accuracy and predictive power.
4.4.1. Pricing Accuracy.
For each portfolio $n$, the return-forecasting regression (Equation 26) yields a pricing error intercept $\alpha_{n}$. We summarize these errors with root‑mean‑squared and mean‑absolute measure (Lettau, 2024):
(27) ${\text{RMSE}_{\alpha}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}\hat{\alpha}_{n}^{2}}},\qquad{\text{MAE}_{\alpha}=\frac{1}{N}\sum_{n=1}^{N}\left|\hat{\alpha}_{n}\right|}.$
Lower values indicate that the selected factors account for a larger share of the cross‑sectional return variation.
4.4.2. Predictive Power.
@@ 4.4. Asset Pricing Evaluation Metrics
Following (Gu et al., 2020), we measure the gap between predicted and realized cross‑sectional rankings with the mean‑absolute rank error:
(28) $\text{MAE‑Rank}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{N}\sum_{n=1}^{N}\left|\operatorname{rank}\left(r_{n,t+1}\right)-\operatorname{rank}\left(\hat{r}_{n,t+1}\right)\right|.$
A smaller MAE-Rank value indicates fewer mis‑ordered portfolios.
The Information Coefficient (IC) (Grinold and Kahn, 2000) captures the linear association between predictions and outcomes:
(29) $\text{IC}=\frac{1}{T}\sum_{t=1}^{T}\operatorname{corr}\left(\hat{\mathbf{r}}_{t+1},\,\mathbf{r}_{t+1}\right),$
@@
To turn ranking skill into investable profits, each month we form a Top-minus-Bottom (T-B) portfolio by longing the decile of firms with the highest predicted excess returns $\hat{r}_{n,t+1}$ and shorting the decile with the lowest (Fama and French, 1993; Carhart, 1997). The resulting excess return is:
(30) $r^{\text{T-B}}_{t+1}=\frac{1}{N_{\text{top}}}\sum_{i\in\text{Top}}r_{i,t+1}-\frac{1}{N_{\text{bot}}}\sum_{i\in\text{Bottom}}r_{i,t+1}.$
where $N_{\text{top}}$ and $N_{\text{bot}}$ are the numbers of firms in the two deciles. We summarize the strategy’s risk‑adjusted performance with its annualized Sharpe ratio (T-B Sharpe):
(31) $\text{Sharpe}_{\text{T-B}}=\sqrt{12}\cdot\frac{\bar{r}^{\text{T-B}}}{\sigma^{\text{T-B}}}.$
@@
5. Experiment
To evaluate the effectiveness of our imputation framework, we adopt a two‑step evaluation that assesses both the statistical accuracy of the data imputation and the financial utility that the imputed data provides in asset‑pricing applications.